OpenStudy (anonymous):

How do you find the linearization L(x) of f(x) = 3x^2+5x+12 at x=3? What are the steps for this?

myininaya (myininaya):

We need to find the tangent line of f at x=3

myininaya (myininaya):

\[f'(x)=6x+15\] so to find the slope use f' and then we also know a point on the line \[(3,3(3)^2+5(3)+12=27+15+12=54)\]

myininaya (myininaya):

\[y=mx+b => 54=23(3)+b \text{ solve for b}\]

OpenStudy (anonymous):

Wouldn't it be f'(x) = 6x + 17, when you take the derivitive?

myininaya (myininaya):

oops i put an extra 1 in there lol

myininaya (myininaya):

\[f'(x)=3(2x)+5=6x+5\]

myininaya (myininaya):

\[(x^2)'=2x;(5x)'=5;(constant)'=0\]

OpenStudy (anonymous):

Ok sounds good, also have another quick question, Where did you get the 23 from... on y=mx+b=>54=23(3)+b solve for b

myininaya (myininaya):

well 23 was what i found when i plug into f' you know what i got earlier for f' lol

myininaya (myininaya):

\[f'(3)=6(3)+5=18+5=23\]

myininaya (myininaya):

f'(3) is the slope of the curve at x=3